A candy mixture is created with 2 types of candy, one costing $4 per pound and the other $3.50 per pound. How much of each type is needed for a 5 pound box that costs $18?

Answer by checkley77(12569)

4x+3.50(5-x)=18
4x+17.5-3.5x=18
4x-3.5x=18-17.5
.5x=.5
x=.5/.5
x=1 pound of the $4.00 candy.
5-1=4 pounds of the $3.50 candy.
Proof:
4*1+3.5*4=18
4+14=18
18=18